# TacSL: A Library for Visuotactile Sensor Simulation and Learning

### v1

@@ Abstract
For both humans and robots, the sense of touch, known as tactile sensing, is critical for performing contact-rich manipulation tasks. Three key challenges in robotic tactile sensing are 1) interpreting sensor signals, 2) generating sensor signals in novel scenarios, and 3) learning sensor-based policies. For visuotactile sensors, interpretation has been facilitated by their close relationship with vision sensors (e.g., RGB cameras). However, generation is still difficult, as visuotactile sensors typically involve contact, deformation, illumination, and imaging, all of which are expensive to simulate; in turn, policy learning has been challenging, as simulation cannot be leveraged for large-scale data collection. We present TacSL (taxel), a library for GPU-based visuotactile sensor simulation and learning. TacSL can be used to simulate visuotactile images and extract contact-force distributions over 200×200$\times$200 × faster than the prior state-of-the-art, all within the widely-used Isaac Gym simulator. Furthermore, TacSL provides a learning toolkit containing multiple sensor models, contact-intensive training environments, and online/offline algorithms that can facilitate policy learning for sim-to-real applications. On the algorithmic side, we introduce a novel online reinforcement-learning algorithm called asymmetric actor-critic distillation (AACD), designed to effectively and efficiently learn tactile-based policies in simulation that can transfer to the real world. Finally, we demonstrate the utility of our library and algorithms by evaluating the benefits of distillation and multimodal sensing for contact-rich manip ulation tasks, and most critically, performing sim-to-real transfer. Supplementary videos and results are at https://iakinola23.github.io/tacsl/.
@@ I Introduction
  • A fast, highly-parallelizable simulation module for visuotactile sensors. Unlike previous works, the module is fully integrated into a widely-used general-purpose robotics simulator, Isaac Gym [13], and provides both simulated RGB images and normal and shear force fields.
@@ II-A Tactile Simulation
For visuotactile sensors, process 1 (contact interaction) involves the indentation and deformation of elastomeric membranes. Gold-standard simulation of elastomeric deformation is typically achieved via the finite element method (FEM) [26]. However, scientific FEM simulators can take exceptionally long to execute [27, 17], less-accurate robotics implementations still execute slower than real-time [28, 29, 30, 17, 31], and faster neural approximations require a large amount of domain-relevant training data [32, 33].
@@ III-A Contact Simulation
This section describes how dynamic contact effects are handled in TacSL. We 1) outline our general dynamics solver procedure, 2) explain the governing analytical equations of our soft contact constraints, and 3) address how these equations are solved in a numerically-stable manner. We implement the necessary contact solver code inside the physics engine and configure the collision geometry and contact parameters from the TacSL side.
Dynamics Solver: Our physics solver operates on a discrete-time formulation with a discretization interval $\Delta t$. Given the position $p$ and velocity $v$ of a body at the current time-step, the solver computes the position $p^{+}$ and velocity $v^{+}$ at the next time step using a semi-implicit Euler integration scheme:
@@
We specifically use a Gauss-Seidel-style solver for its simplicity and fast convergence. The solver subdivides the frame duration into a configurable number of smaller timesteps and applies a sequential impulse strategy for each of these substeps. Each constraint (e.g., a collision) computes an impulse to reduce the constraint error and applies it to the system by changing $\Delta v$ of the involved bodies. The procedure for one time frame is given in Algorithm 1.
Algorithm 1 Dynamics Solver Step
@@
4:         Compute the impulse $\lambda$ that is to be applied over this substep based on the most up-to-date estimate of the positions and velocities.
5:         Update the body velocities based on the impulse.
@@
7:     Integrate body position based on the velocity from the constraint solve over the timestep $\Delta t/N$.
@@ III-C Normal/Shear Force-Field Computation
Prior to simulation, we compute the signed distance field (SDF) of the contacting object, as well as the gradient of the field, as described in [49, 50]. At each timestep, to get the interpenetration distance $d$, we query the signed distance field of the contacting object at the tactile points. To get the interpenetration velocity $\dot{d}$, we use chain rule to rewrite $\dot{d}=(\mathbf{\nabla}{d})^{T}\mathbf{\dot{x}}$, where $\mathbf{\dot{x}}$ is the relative velocity at the contact point. To get $\mathbf{n}$, we use the relation $\mathbf{\nabla}{d}=\mathbf{n}$.
@@ IV-A Policy Distillation
TacSL implements two standard approaches to policy distillation. First, offline distillation, also referred to as behavior cloning (BC), learns to mimic actions from a fixed dataset of simulated experience. The simplicity of BC enables computationally-cheap and fast implementations; however, its performance is limited to the size and quality of the offline dataset. Second, TacSL provides an online distillation approach, also known as DAgger[52], where the student policy is increasingly used to collect more experience as training proceeds. For completeness, we describe our application of the approach in detail in Algorithm 2 and Fig. 4.
@@ IV-B2 Asymmetric Actor-Critic Distillation (AACD)
Inspired by the teacher-student framework described earlier, we present a novel asymmetric actor-critic distillation algorithm (AACD) that addresses this exploration challenge by leveraging a pretrained critic from a low-dimensional agent to guide the learning process of a new high-dimensional agent. AACD trains a high-dimensional policy in two steps (see Fig. 5). First, a randomly-initialized RL actor and critic with low-dimensional observation input are trained from scratch using low-dimensional privileged state information. Next, a randomly-initialized actor with high-dimensional observation input and the pretrained low-dimensional critic are trained and fine-tuned, respectively, to optimize RL policy objectives. AACD effectively addresses exploration challenges in high-dimensional policy learning by leveraging prior knowledge of the pretrained critic, thereby providing effective guidance for learning the high-dimensional policy.
@@ V-A Physics Parameter Randomization
To account for the variation in elastomer compliance across sensors and potential changes over time, we randomize the parameters of the soft contact model of our simulator. The stiffness and damping parameters of our compliant contact model determine the level of softness and velocity decay rate, respectively, of the soft contacts in simulation. These parameters are empirically tuned to reasonable values and are then randomized during policy training (see randomization details in Appendix B). This randomization ensures that the agent can adapt to uncertainties in the model parameters and variations in the real-world environment.
@@ VI-A Tactile Simulation Speed
Table I shows that the effective speed of our simulator increases as we increase the number of simulated parallel environments before the GPU gets saturated. Our simulator runs at $1631$ frames per second (FPS) with $512$ parallel environments on an NVIDIA RTX 3090, compared to the baseline (Taxim [18]) that runs at $7.28$ FPS on a single core of an AMD Threadripper 1950X processor. This represents a more than 200×200$\times$200 × speed-up by TacSL over a baseline with comparable tactile-image quality. This significant performance improvement is attributed to TacSL full utilization of the GPU-based parallel processing capabilities available in Isaac Gym, with most computations performed on the GPU and minimal data transfer overhead between the GPU and CPU.
For tactile shear force fields, we compare the simulation speed at two tactile force field resolutions, $10\times 10$ and $100\times 100$. The $10\times 10$ resolution captures the typical tactile shear resolution used by most real-world visuotactile sensors (e.g. GelSight) use, providing a speed evaluation for practical scenarios. On the other side, $100\times 100$ represents a stress test of the simulation’s speed on a high-resolution tactile sensing field. Table II shows that for low-resolution tactile field ($10\times 10$), the speed of our simulator linearly scales up with the number of environments. When we use $32768$ parallel environments for the ball-rolling task, our simulator achieves $1541043$ FPS which is 428×428$\times$428 × speedup over the CPU baseline [16]. For the high-resolution stress test case ($100\times 100$), the speed of our simulation gets saturated when we increase the number of parallel environments to $4096$. With $4096$ parallel environments, our simulator is able to simulate $100\times 100$ tactile force field in $103493$ FPS, which is still 46×46$\times$46 × speedup over the baseline.
@@ VI-B1 Tasks
  • Peg Insertion. Here the robot has to insert a cylindrical rod into a cylindrical socket. Similar to the Peg Placement task, the rod starts at a random position and orientation within the robot’s gripper.
@@ Acknowledgments
We would like to thank our colleagues at NVIDIA for their invaluable assistance and feedback throughout this work. Special acknowledgment to Michael Noseworthy, Bingjie Tang, Bowen Wen, Karl Van Wyk, Ankur Handa and Fabio Ramos for engaging in deep conversation and providing insightful feedback. We deeply thank Philipp Reist for his insights on the physics solver and detailed feedback on the paper draft. We appreciate the assistance of Tobias Widmer and Milad Rakhsha in the SDF implementation. Our thanks also go to Viktor Makoviychuk and Kelly Guo for their responsiveness to our inquiries about Isaac Gym. We thank Ajay Mandlekar for assistance in setting up behavior cloning during early prototyping, and Alperen Degirmenci for help in creating visualizations and figures for the paper. Valuable feedback on the paper draft was provided by Yu-Wei Chao. We express our thanks to Kimo Johnson for important GelSight hardware support.
Appendix A Contact Model Calibration
@@
Appendix B Training Details
@@
Reward function: Adapted from [49], the reward function is given as:
@@
  • $r_{\text{keypoint}}$ is distance between keypoints centered on the peg and keypoints of its target pose on the placement pad/socket.
@@
B-A Policy Learning with different assets.
@@
Appendix C Tactile Force Field Edge-Test
@@
Appendix D Tactile Image Visualization
@@
Appendix E Additional Tactile Visualizations
We present additional visualizations (Figure LABEL:fig) of both tactile images and tactile force fields for the GelSight R1.5 sensor interacting with a complex bolt mesh, which comprises 30k faces and 26k vertices, at various poses.
Figure 17: Additional visualizations of tactile readings (tactile image and shear force field) from the GelSight 1.5 sensor as the bolt mesh interacts with the sensor. The bolt, positioned at varying poses, applies a downward force and rotates in the anticlockwise direction.
Appendix F Biography Section
[Uncaptioned image] Iretiayo Akinola Iretiayo Akinola received a PhD from Columbia University’s Computer Science Department in 2021, where his research centered on equipping robots with multi-modal (visual and tactile) perception for reactive object grasping and manipulation. Prior to that, he earned his M.S. and BSc. degrees in Electrical Engineering from Stanford University and Obafemi Awolowo University, respectively. His research interests include multi-modal robot learning, sim-to-real transfer, and human-in-the-loop robot learning.
[Uncaptioned image] Jie Xu Jie Xu received the B.Eng. degree from Department of Computer Science and Technology at Tsinghua University in 2016 and Ph.D. degree in Computer Science from Massachusetts Institute of Technology in 2022. He is currently a Research Scientist at Nvidia Research. His research interests include Robotics Control, Reinforcement Learning, Simulation, Robot Co-Design and Sim-to-Real.
[Uncaptioned image] Jan Carius Jan Carius is currently working in the domain of physics simulations and parallel computing. He received a PhD from ETH Zurich in 2021 where his research evolved around motion planning and control algorithms for legged locomotion based on optimal control and machine learning. Formerly he received BSc. and MSc. degrees in mechanical engineering from ETH Zurich.
[Uncaptioned image] Dieter Fox Dieter Fox received his PhD degree from the University of Bonn, Germany. He is a professor in the Allen School of Computer Science & Engineering at the University of Washington, where he heads the UW Robotics and State Estimation Lab. He is also Senior Director of Robotics Research at NVIDIA. His research is in robotics and artificial intelligence, with a focus on learning and estimation applied to problems such as robot manipulation, planning, language grounding, and activity recognition. He has published more than 300 technical papers and is co-author of the textbook “Probabilistic Robotics”. Dieter is a Fellow of the IEEE, ACM, and AAAI, and recipient of the IEEE RAS Pioneer Award and the IJCAI John McCarthy Award.
[Uncaptioned image] Yashraj Narang Yashraj Narang received a B.S. from Stanford University in 2011, an S.M. from the Massachusetts Institute of Technology in 2013, and a Ph.D. from Harvard University in 2018. He joined the NVIDIA Seattle Robotics Lab as a research scientist in 2018 and has led the Simulation and Behavior Generation group since 2023. His current research focuses on leveraging physics-based simulation for robot learning in structured and unstructured environments.

### v2

@@ Abstract
For both humans and robots, the sense of touch, known as tactile sensing, is critical for performing contact-rich manipulation tasks. Three key challenges in robotic tactile sensing are 1) interpreting sensor signals, 2) generating sensor signals in novel scenarios, and 3) learning sensor-based policies. For visuotactile sensors, interpretation has been facilitated by their close relationship with vision sensors (e.g., RGB cameras). However, generation is still difficult, as visuotactile sensors typically involve contact, deformation, illumination, and imaging, all of which are expensive to simulate; in turn, policy learning has been challenging, as simulation cannot be leveraged for large-scale data collection. We present TacSL (taxel), a library for GPU-based visuotactile sensor simulation and learning. TacSL can be used to simulate visuotactile images and extract contact-force distributions over 200×200$\times$200 × faster than the prior state-of-the-art, all within the widely-used Isaac Simulator. Furthermore, TacSL provides a learning toolkit containing multiple sensor models, contact-intensive training environments, and online/offline algorithms that can facilitate policy learning for sim-to-real applications. On the algorithmic side, we introduce a novel online reinforcement-learning algorithm called asymmetric actor-critic distillation (AACD), designed to effectively and efficiently learn tactile-based policies in simulation that can transfer to the real world. Finally, we demonstrate the utility of our library and algorithms by evaluating the benefits of distillation and multimodal sensing for contact-rich manipulation tasks, and most critically, performing sim-to-real transfer. Supplementary videos and results are at https://iakinola23.github.io/tacsl/.
@@ I Introduction
  • A fast, highly-parallelizable simulation module for visuotactile sensors. Unlike previous works, the module is fully integrated into a widely-used general-purpose robotics simulator, Isaac Simulator [13], and provides both simulated RGB images and normal and shear force fields.
@@ II-A Tactile Simulation
For visuotactile sensors, process 1 (contact interaction) involves the indentation and deformation of elastomeric membranes. Gold-standard simulation of elastomeric deformation is typically achieved via the finite element method (FEM) [26]. However, scientific FEM simulators can take exceptionally long to execute [27, 17], less-accurate robotics implementations still execute slower than real-time [28, 29, 30, 17, 31], and faster neural approximations require a large amount of domain-relevant training data [32, 33]. A recent work [34] developed a FEM-based physics engine for simulating the tactile sensor deformation accurately and efficiently. However, the authors identified several limitations, including low simulation speed when handling complex rigid bodies and the lack of integration with a general-purpose simulator.
@@ III-A Contact Simulation
This section describes how dynamic contact effects are handled in TacSL. We 1) outline our general dynamics solver procedure, 2) explain the governing analytical equations of our soft contact constraints, and 3) address how these equations are solved in a numerically-stable manner. The contact constraint equations derived here are implemented within the PhysX SDK [46] physics engine, whose built-in solver procedure (outlined below) ensures constraint-consistent dynamics. Collision geometry and contact parameters are configured on the TacSL side.
Dynamics Solver: The PhysX dynamics solver operates on a discrete-time formulation with a discretization interval $\Delta t$. Given the position $p$ and velocity $v$ of a body at the current time-step, the solver computes the position $p^{+}$ and velocity $v^{+}$ at the next time step using a semi-implicit Euler integration scheme:
@@
We use a Gauss-Seidel-style solver due to its simplicity and fast convergence, based on [47]. The solver subdivides the frame duration into a configurable number of smaller timesteps and applies a sequential impulse strategy for each of these substeps. Each constraint (e.g., a collision) computes an impulse to reduce the constraint error and applies it to the system by changing $\Delta v$ of the involved bodies. The procedure for one time frame is given in Algorithm 1, with further details of the dynamics solver provided in Appendix A.
Algorithm 1 Dynamics Solver Step (Gauss-Seidel)
@@
4:         Compute the impulse $\lambda$ that is to be applied over this substep based on the most up-to-date estimate of the positions and velocities, ignoring any other constraints.
5:         Update the body velocities based on the computed impulse.
@@
7:     Integrate body positions based on the updated velocities over the timestep $\Delta t/N$.
@@ III-C Normal/Shear Force-Field Computation
Prior to simulation, we compute the signed distance field (SDF) of the contacting object as described in [52, 53]. At each timestep, the interpenetration distance $d$ is obtained by querying the SDF of the contacting object at the tactile points. To get the interpenetration velocity $\dot{d}$, we use chain rule to rewrite $\dot{d}=(\mathbf{\nabla}{d})^{T}\mathbf{\dot{x}}$, where $\mathbf{\dot{x}}$ is the relative velocity at the contact point. The contact normal $\mathbf{n}$ is defined as the gradient of the SDF values: $\mathbf{\nabla}{d}=\mathbf{n}$. The SDF gradient is calculated using finite differencing. Notably, both SDF queries and gradient computations are computationally efficient and easily parallelizable.
@@ IV-A Policy Distillation
Two standard approaches to policy distillation are implemented in TacSL. First, offline distillation, also referred to as behavior cloning (BC), learns to mimic actions from a fixed dataset of simulated experience. The simplicity of BC enables computationally-cheap and fast implementations; however, its performance is limited to the size and quality of the offline dataset. Second, TacSL provides an online distillation approach, also known as DAgger[55], where the student policy is increasingly used to collect more experience as training proceeds. For completeness, we describe our application of the approach in detail in Algorithm 2 and Fig. 4.
@@ IV-B2 Asymmetric Actor-Critic Distillation (AACD)
Algorithm 3 Asymmetric Actor-Critic Distillation
1:initial low-dim policy parameters $\theta_{s}^{0}$, low-dim value function parameters $\phi_{s}^{0}$, and high-dim policy parameters $\theta_{o}^{0}$. All parameters are randomly initialized.
2:Stage 1: Train low-dim policy $\pi_{\theta_{s}}(s)$ and value function $V_{\phi_{s}}(s)$
3:for $n=0$ to $N-1$ do
4:     Collect rollout data with $\pi_{\theta_{s}^{n}}(s)$
5:     Update $\theta_{s}^{n+1}$ using PPO objective and $V_{\phi_{s}^{n}}(s)$
6:     Update $\phi_{s}^{n+1}$ using Bellman loss
7:end for
8:Stage 2: Train high-dim policy $\pi_{\theta_{o}}(o)$ and fine-tune $V_{\phi_{s}}(s)$
9:for $m=0$ to $M-1$ do
10:     Collect rollout data with $\pi_{\theta_{o}^{m}}(o)$
11:     Update $\theta_{o}^{m+1}$ using PPO objective and $V_{\phi_{s}^{N+m}}(s)$
12:     Update $\phi_{s}^{N+m+1}$ using Bellman loss
13:end for
14:Evaluate $\pi_{\theta_{o}^{M}}$triangleright Obtain success rate
Inspired by the teacher-student framework described earlier, we present a novel asymmetric actor-critic distillation algorithm (AACD) that addresses this exploration challenge by leveraging a pretrained critic from a low-dimensional agent to guide the learning process of a new high-dimensional agent. The low-dimensional agent employs a privileged state-based policy, while the high-dimensional agent operates with image-based observations. AACD trains a high-dimensional policy in two steps (see Fig. 5 and Algorithm 3). First, a randomly-initialized RL actor and critic with low-dimensional observation input are trained from scratch using low-dimensional privileged state information. Next, a randomly-initialized actor with high-dimensional observation input and the pretrained low-dimensional critic are trained and fine-tuned, respectively, to optimize RL policy objectives. AACD effectively addresses exploration challenges in high-dimensional policy learning by leveraging prior knowledge of the pretrained critic, thereby providing effective guidance for learning the high-dimensional policy.
@@ V-A Physics Parameter Randomization
To account for the variation in elastomer compliance across sensors and potential changes over time, we randomize the parameters of the soft contact model of our simulator. The stiffness and damping parameters of our compliant contact model determine the level of softness and velocity decay rate, respectively, of the soft contacts in simulation. These parameters are empirically tuned to reasonable values and are then randomized during policy training (see randomization details in Appendix C). This randomization ensures that the agent can adapt to uncertainties in the model parameters and variations in the real-world environment.
@@ VI-A Tactile Simulation Speed
Table I shows that the effective speed of our simulator increases as we increase the number of simulated parallel environments before the GPU gets saturated. Our simulator runs at $1631$ frames per second (FPS) with $512$ parallel environments on an NVIDIA RTX 3090, compared to the baseline (Taxim [18]) that runs at $7.28$ FPS on a single core of an AMD Threadripper 1950X processor. This represents a more than 200×200$\times$200 × speed-up by TacSL over a baseline with comparable tactile-image quality. This significant performance improvement is attributed to TacSL full utilization of the GPU-based parallel processing capabilities available in Isaac Simulator, with most computations performed on the GPU and minimal data transfer overhead between the GPU and CPU.
For tactile shear force fields, we compare the simulation speed at two tactile force field resolutions, $10\times 10$ and $100\times 100$. The $10\times 10$ resolution captures the typical tactile shear resolution used by most real-world visuotactile sensors (e.g. GelSight), providing a speed evaluation for practical scenarios. On the other side, $100\times 100$ represents a stress test of the simulation’s speed on a high-resolution tactile sensing field. Table II shows that for low-resolution tactile field ($10\times 10$), the speed of our simulator linearly scales up with the number of environments. When we use $32768$ parallel environments for the ball-rolling task, our simulator achieves $1541043$ FPS which is 428×428$\times$428 × speedup over the CPU baseline [16]. For the high-resolution stress test case ($100\times 100$), the speed of our simulation gets saturated when we increase the number of parallel environments to $4096$. With $4096$ parallel environments, our simulator is able to simulate $100\times 100$ tactile force field in $103493$ FPS, which is still 46×46$\times$46 × speedup over the baseline.
@@ VI-B1 Tasks
  • Peg Insertion. Here the robot has to insert a cylindrical rod into a cylindrical socket[52, 61]. Similar to the Peg Placement task, the rod starts at a random position and orientation within the robot’s gripper.
@@ Acknowledgments
We would like to thank our colleagues at NVIDIA for their invaluable assistance and feedback throughout this work. Special acknowledgment to Michael Noseworthy, Bingjie Tang, Bowen Wen, Karl Van Wyk, Ankur Handa and Fabio Ramos for engaging in deep conversation and providing insightful feedback. We deeply thank Philipp Reist for his insights on the physics solver and detailed feedback on the paper draft. We appreciate the assistance of Tobias Widmer and Milad Rakhsha in the SDF implementation. Our thanks also go to Viktor Makoviychuk and Kelly Guo for their responsiveness to our inquiries about Isaac Simulator. We thank Ajay Mandlekar for assistance in setting up behavior cloning during early prototyping, and Alperen Degirmenci for help in creating visualizations and figures for the paper. Valuable feedback on the paper draft was provided by Yu-Wei Chao. We express our thanks to Kimo Johnson for important GelSight hardware support.
Appendix A Dynamics Solver Details
TacSL employs the Temporal Gauss-Seidel (TGS) algorithm, implemented in the NVIDIA PhysX SDK [46], as its dynamics solver. The TGS solver is an advanced iterative method designed for efficiently and robustly resolving constraints in systems with complex interactions, such as collisions, friction, and joints. The solver divides each frame duration into $N$ substeps, reducing the integration timestep to $\Delta t/N$. During each substep, constraints are processed sequentially.
Constraint Resolution
For each constraint, such as a collision or joint limit, the TGS solver calculates an impulse $\lambda$ to minimize the constraint error. This impulse is computed using the constraint’s gradient and its compliance (if any), incorporating temporal stabilization terms specific to TGS. These stabilization terms address errors from discrete timesteps by incorporating positional corrections accumulated over previous time steps into the velocity update [47], thereby preventing error accumulation over time. The computed impulse is then applied to the bodies affected by the constraint by adjusting their velocities $\Delta v$. The sequential nature of the Gauss-Seidel approach ensures that the most recent updates to positions and velocities are used for subsequent constraints, which promotes convergence and enhances stability in tightly coupled systems.
Integration
After processing all constraints for a substep, the solver integrates the bodies’ positions using the updated velocities. This substep integration maintains consistency between positions and velocities throughout the simulation. The combination of substepping and TGS’s stabilization minimizes errors from discrete timesteps and improves the handling of dynamic interactions.
For further implementation details, including the source code, we refer the reader to the open-source PhysX SDK repository [46].
Appendix B Contact Model Calibration
@@
Appendix C Training Details
@@
Reward function: Adapted from [52, 67], the reward function is given as:
@@
  • $r_{\text{keypoint}}$ is the distance between keypoints centered on the peg and keypoints of its target pose on the placement pad/socket, passed through an exponential function.
@@
C-A Policy Learning with different assets.
@@
Appendix D Tactile Force Field Edge-Test
@@
Appendix E Tactile force-field notation
The tactile force-field computation is a separate post-processing step after each dynamics simulation step. In principle, both the parameter $\epsilon$, used in the formulation of the dynamics solver of the full simulator (Section III-A), and the parameter $d$, used in the penalty-based computation of tactile force-field (Section III-C), refer to the same quantity– penetration depth. However, they refer to the interpenetration depth of different categories of points. The dynamics solver generates contact points based on objects in the scene when resolving collision constraints, while the tactile field computation is done on predefined tactile points which are a sampled grid on the tactile sensor and correspond to the markers that are embedded in the physical sensor.
Appendix F Tactile Image Visualization
@@
Appendix G Additional Tactile Visualizations
We present additional visualizations (Figure 17) of both tactile images and tactile force fields for the GelSight R1.5 sensor interacting with a complex bolt mesh, which comprises 30k faces and 26k vertices, at various poses.
Figure 17: Additional visualizations of tactile readings (tactile image and shear force field) from the GelSight 1.5 sensor as the bolt mesh (30k faces and 26k vertices) interacts with the sensor. The bolt, positioned at varying poses, applies a downward force and rotates in the anticlockwise direction.